OpenStudy (anonymous):

d/dx x^4x is there a proof?

OpenStudy (amistre64):

a proof for what?

OpenStudy (amistre64):

even if its ^(4x) the power rule and chain rules apply

OpenStudy (anonymous):

d/dx a^x=a^xlna

OpenStudy (anonymous):

Yes just apply the chain rule.

OpenStudy (anonymous):

can you work it out for me?

OpenStudy (anonymous):

\[(f(g(x))' = f'(g(x))g'(x)\]

OpenStudy (amistre64):

f(x) = x^(g(x)) g(x) = 4(h(x)) h(x) = x

OpenStudy (anonymous):

huh how come it becomes a function...?

OpenStudy (anonymous):

composite function..

OpenStudy (amistre64):

just by the nature of what a function is

OpenStudy (amistre64):

you can rename the parts of an expression/equation to match seperate function

OpenStudy (amistre64):

sqrt(5x) for example is: sqrt(...) 5(...) x

OpenStudy (amistre64):

it helps to see the chain rule in action

OpenStudy (anonymous):

i cant see a chain rule..

OpenStudy (anonymous):

whats chain rule

OpenStudy (zarkon):

you want this ... \[\frac{d}{dx}x^{4x}\]

OpenStudy (amistre64):

f(x) = x^(g(x)) f'(x) = g(x) x^(g(x)-1) g(x) = 4(h(x)) g'(x) = 4(h'(x)) h(x) = x h'(x) = 1

OpenStudy (zarkon):

that is not true

OpenStudy (amistre64):

the chain rule is akin to peeling an onion ... and deriving each section

OpenStudy (zarkon):

this .... f(x) = x^(g(x)) f'(x) = g(x) x^(g(x)-1) is not true

OpenStudy (anonymous):

then how to solve?

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=d%2Fdx+x^%284x%29 has steps if im wrong

OpenStudy (zarkon):

\[\frac{d}{dx}x^{4x}=\frac{d}{dx}e^{4x\ln(x)}\]

OpenStudy (zarkon):

now use chain and product rule

OpenStudy (anonymous):

Maybe he/she want definition of derivative

OpenStudy (amistre64):

im open to being mistaken at times ;)

OpenStudy (anonymous):

zarkon how did you arrive at that step

OpenStudy (zarkon):

\[\frac{d}{dx}x^{4x}=\frac{d}{dx}e^{4x\ln(x)}\] \[=e^{4x\ln(x)}\left(\frac{4x}{x}+4\ln(x)\right)\] \[=x^{4x}\left(4+4\ln(x)\right)\]

OpenStudy (anonymous):

how did the first step come about

OpenStudy (zarkon):

\[y=x^{4x}\] \[\ln(y)=\ln(x^{4x})\] \[\ln(y)=4x\ln(x)\] \[y=e^{4x\ln(x)}\]

OpenStudy (zarkon):

ok?

OpenStudy (anonymous):

OK! thanks=)

OpenStudy (anonymous):

so whenever you see that you ln both sides?

OpenStudy (zarkon):

\[\frac{d}{dx}h(x)^{g(x)}\] do this \[\frac{d}{dx}e^{g(x)\ln(h(x))}\]

OpenStudy (zarkon):

just never use the power rule. that only applies when the exponent is a constant with respect to the variable you are differentiating with.

OpenStudy (anonymous):

is that memory work?

OpenStudy (zarkon):

?

OpenStudy (anonymous):

that step you type when you first see the question

OpenStudy (zarkon):

that's what I usually do

OpenStudy (anonymous):

how do u see it

OpenStudy (zarkon):

when I see a function of x raised to a function of x I always rewrite it in terms of exponential and log functions

OpenStudy (zarkon):

that is...if i plan to differentiate